OpenStudy (anonymous):

what are three different representations of (4, 30) on a polar graph?

OpenStudy (anonymous):

\[x = r \cos \theta, y = r \sin \theta\] Here x = 4, y = 30.

OpenStudy (anonymous):

you need to solve for r and theta

OpenStudy (phi):

if you add a multiple of 360º to the angle, you get the same point. So there are lots of different ways to write the same number in polar coordinates.

OpenStudy (anonymous):

ohhh i get it so it's more like this (x=rcosø), y=rsinø) they're like coordinates?

OpenStudy (phi):

if (4,30) represents r and theta, then add 360 to the angle if (4,30) is rectangular coordinates, you need to change to polar form.

OpenStudy (anonymous):

the polar coordinate would be (r, theta)

OpenStudy (anonymous):

no no i get that they are polar coordinates but all I do is add 360? won't that give me the same number?

OpenStudy (anonymous):

i mean unless I go farther. Is one of them (4, 390)?

OpenStudy (phi):

yes (4,390) is another "representation" for the point (4,30)

OpenStudy (anonymous):

okay! thanks!

OpenStudy (phi):

you can also add 720 degrees or subtract 360 or, slightly more complicated, add 180º and change the sign of r: (-4, 210º) for example. But it is best not to use negative r values. (you can always make them positive by adding or subtracting 180º from the angle)

OpenStudy (anonymous):

okay awesome! I think I get it now!

OpenStudy (phi):

yw